OpenStudy (anonymous):

Lyle has a ten-cup box of cereal. If one serving is 3/4 cup, how many servings are in the box?

OpenStudy (anonymous):

You divide the total by the servings. So its: \[10 / (3/4) = 10 (4/3) = 40 /3 = 13.3\]